Please help me solve this equation: the length of a rectangle is 4 centimeters more than twice its width. If the area of the recatangle is 160 cm, find the dimensions.
Answer by Alan3354(69443)

Please help me solve this equation: the length of a rectangle is 4 centimeters more than twice its width. If the area of the recatangle is 160 cm, find the dimensions.
------------------
You didn't post an equation.
---
L = 2W + 4
L*W = 160
Sub for L
W*(2W + 4) = 160
W^2 + 2W - 80 = 0
W = 8